# On Exact Space-Depth Trade-Offs in Multi-Controlled Toffoli Decomposition

### v1

@@ I Introduction
Quantum gates are the fundamental building blocks of quantum circuits. Unlike classical gates, quantum gates are inherently reversible and are mathematically represented by unitary matrices. The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations, reversible computing, and oracle constructions. However, the Toffoli gate’s high resource demands, particularly in terms of T-Count, T-Depth, and ancilla qubits, can greatly influence the efficiency of fault-tolerant quantum circuits. Consequently, optimizing its implementation is crucial for reducing computational overhead, minimizing error rates, and enhancing scalability, thereby making it indispensable for practical large-scale quantum computations.
An $n$-controlled Toffoli is an $(n+1)$-qubit gate, where $n$ many qubits control the outcome of a single target qubit. Like ($2$-controlled) Toffoli, Multi Controlled Toffoli (MCT) also has a huge applications in quantum arithmetic, error correction, and implementation of quantum algorithms. For example, in Grover’s Search, implementing the $n$-bit AND function requires an $n$-controlled Toffoli ($n$-MCT) gate. As MCT cannot be implemented in its original form (so far), it needs to be decomposed into Clifford + Toffoli, and eventually to the Clifford + T gate set. Throughout the paper, By Toffoli, we refer to $n$-controlled Toffoli with $n=2$.
It is well known that decomposing/ designing larger circuits with smaller components requires additional qubits, known as ancilla qubits, to reduce depth. The important task for optimizing the number of gates, the depth and the number of additional qubits has received serious attention for a long time and one may refer to the seminal work of Moore and Nilsson moore01 two decades back. A more recent and comprehensive discussion is available in jiang20 and we like to quote from that work “Can we characterize the relationship between the number of ancilla and the possible optimal depth?” The paper jiang20 is concerned with CNOT, whereas in this paper we concentrate on Toffoli. However, the fundamental motivation of quantum circuit synthesis revolves around this very question.
@@
Before proceeding further let us explain the concept of conditionally clean ancilla that will be repeatedly referred in this work. Generally, ancilla qubits are of two types, namely the clean ancilla, which is initialized to |$\ket{0}$ at the beginning of the circuit, and the dirty ancilla, which is initialized to some unknown quantum state at the beginning. The standard practice is to re-initialize the ancilla qubit to its initial state at the end of the computation. Conditionally clean ancilla nie24 ; khattar24 are the set of working (control) qubits, derived form the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation. Although, earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancilla derived from the clean ones only. Very recently, the conditionally clean ancilla technique has gained significant attention, including its applications in quantum adders remaud25 .
The resource optimization metrics under consideration are the Toffoli count , Toffoli depth and Ancilla count. The Toffoli count and Toffoli depth are further refined with the T-Count and T-Depth as one may refer to Table 1.
@@ I.1 Organization and Contributions
In Section II, we proceed with the preliminaries. Section III is a warm-up section where we first present the existing results on the Clifford $+$ T decomposition of ($2$-controlled) Toffoli and summarize the results in terms of T-Count, T-Depth, and ancilla requirements in Table 1. Additionally, we also explain the recent developments of MCT decompositions describing the existing best results in terms of Toffoli count, Toffoli depth, and ancilla, which can be subsequently decomposed into T-Count and T-Depth, as summarized in Table 2. Given that there are several developments in very recent times, this section provide a holistic view of the existing results. Based on this, we present our contributions.
In section IV, we explore the trade-off between the (clean) ancilla and the Toffoli depth using the existing techniques related to conditionally clean ones. In Section IV.1, we take a different look in viewing the MCT circuit decomposition using conditionally clean ancilla technique of khattar24 , and enumerate their exact Toffoli count. Following the trade-off, we show the reduction in Toffoli depth, and therefore in T-Depth (Construction 1, Example 1) compared to the recent work of Khattar and Gidney khattar24 , by introducing additional clean ancilla into the circuit, while keeping the Toffoli count constant. These results are shown in Section IV.2. Then, in Section IV.3, we identify the limitation of this technique in Theorem 3, showing that this direction cannot reduce the Toffoli depth to $\lceil\log_{2}n\rceil$, though it is of order $\mathcal{O}(\log_{2}n)$.
In Section V, we prove within a general framework that the exact Toffoli depth in the ($2$-controlled) Toffoli decomposition of an $n$-MCT cannot be less than $\lceil\log_{2}n\rceil$, regardless of the number of ancilla qubits used. In fact, this lower bound can be achieved for T-Depth as well, by the construction of jaques19 . That is, using the technique of jaques19 , we can obtain an $n$-MCT by further decomposing into the Clifford + T gate set with an exact T-Depth of $\lceil\log_{2}n\rceil$ too, using $22n-$ 2 ancilla, and a T-Count of $4(n-1)$. Moreover, using the logical-AND circuit by Gidney gidney18 , the ancilla count can be reduced to $n-2$ keeping the T-Count constant. However, the exact T-Depth in case of gidney18 becomes one more, i.e., $\lceil\log_{2}n\rceil+1$. To highlight our contribution, we are looking at the exact counts and depth instead of their complexity order.
@@ II Preliminaries
However, the Clifford gates alone do not constitute a universal set of quantum gates because certain gates. For example, the T $=\sqrt{\text{S}}$ gate, cannot be arbitrarily approximated using only Clifford operations. Therefore, the Clifford group, when augmented with the T gate, forms a universal quantum gate set for quantum computation.
@@
In quantum computing, the Multi-Controlled Toffoli (MCT) gates play a crucial role in the design of complex quantum algorithms, including error correction codes and arithmetic operations. Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set, requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancillary qubits. These optimizations are essential for practical quantum computation, where resource efficiency is a critical consideration. In this regard, we present a comprehensive overview of the existing benchmarks for Clifford + T decompositions of single and multi-controlled Toffoli gates in the following section.
Let us now describe the idea of conditionally clean ancilla from khattar24 , as described in Figure 2. Given a clean ancilla, a Toffoli gate is implemented targeting the clean ancilla. If the clean ancilla is reversed, it means both the control qubits were |$\ket{1}$; thus, applying X gates on the control qubits will change them to |$\ket{0}$ and consequently can be used as conditionally clean ancilla in the next rounds. Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancilla and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits. Once the information from all the control qubits is accumulated in a few qubits, another (smaller) MCT gate is implemented with these qubits along with the ancilla as the control qubits and the original target as its target, so that if the ancilla was not modified in the first step, the target will also not be modified.
@@
Case II: At least one of $C_{i}=\ket{0}$, for some $i\in[0,4]$. If one of $C_{0}$ or $C_{1}$ is |$\ket{0}$, then ancilla =|$\ket{0}$, and the target will not be flipped in Step III. If $C_{0}=C_{1}=\ket{1}$, but one of $C_{2},C_{3},C_{4}$ is |$\ket{0}$, then ancilla becomes |$\ket{1}$, and after applying X-gate on $C_{0},C_{1}$ will make, $C_{0}=C_{1}=\ket{0}$. If one of $C_{2},C_{3}$ is |$\ket{0}$, then $C_{0}$ and consequently, $C_{1}$ also remains at |$\ket{0}$ state. Similarly, if $C_{4}=\ket{0}$, then also $C_{1}$ remains |$\ket{0}$, and in both the scenario, the target will not be flipped in Step III.
@@ III.2 Multi-Controlled Toffoli Decomposition
In Feb 2024, Nie et al. nie24 first used the notion of conditionally clean ancilla qubits derived from an existing (clean or dirty) ancilla and proposed a novel circuit decomposition (Figure 10) for the $n$-controlled Pauli gates using $\mathcal{O}(n)$ Toffoli gates. The outer layer of their MCT decomposition follows from gidney15 , and the inner layer has been parallelized to obtain an overall Toffoli depth of $\mathcal{O}(\log_{2}n)$, compared to the $\mathcal{O}(\log_{2}n)$ Toffoli depth in gidney15 . Additionally, by improving the design of a quantum incrementer, they developed an MCT circuit with a Toffoli count of $\mathcal{O}(n)$, and a Toffoli depth of $\mathcal{O}(\log^{2}n)$ without requiring any additional ancilla qubits. In this design, although the MCT decomposition does not use any additional ancilla, implementing the quantum incrementer requires one ancilla. Since, our primary focus here is to reduce Toffoli depth while increasing the clean ancilla count, we ignore the zero-ancilla implementation here.
@@
The MCT circuit decomposition proposed by nie24 using a single clean ancilla, has a minimum Toffoli count of $44n+$ 4, and an exact Toffoli depth of $20\log_{2}n$.
@@ IV.1 Exact Enumeration of Toffoli Depth proposed by Khattar and Gidney (2024) in a Different Lens
Using this new approach the $5$-MCT circuit using conditionally clean ancilla technique, presented in Figure 2, can be seen as follows.
@@
where the Toffoli depth for Step I and Step II combined is 3, which can also be verified from the actual circuit from Figure 2. Moreover, a $2$-controlled Toffoli needs to implemented in Step III to modify the target.
@@
In the direction of enumerating exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancilla, we have the following lemma.
@@
From above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using 2 ancilla qubits, as proposed in khattar24 .
@@
where $k\in\mathbb{N}$, with $k\geq 2$. Since, here we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range, and still show that $3\log_{2}n$ is strictly less than the corresponding depths, as follows.
@@
Although the above proposition establishes that the exact Toffoli depth using two clean ancilla, as per khattar24 , is lower bounded by $3\log_{2}n$, this bound is not tight and is closer to $4\log_{2}n$. In the following section, we analyze the trade-off between Toffoli depth and clean ancilla, demonstrating that with additional ancilla, the exact Toffoli depth can be reduced to $2\log_{2}n$.
@@ IV.2 Exact Trade-Off between Toffoli Depth and Clean Ancillae
The schematic diagram of the MCT decomposition using $m=m_{1}+m_{2}$ many ancilla is similar to the diagram we used for the exact Toffoli depth enumeration in Section IV.1.
@@
The Toffoli depth of the MCT decomposition using $m=m_{1}+m_{2}$ many ancilla, can be estimated from the following lemma.
@@
The exact Toffoli depth of the Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancilla following the conditionally clean ancilla technique, varies with $n$ as follows.
@@
The number of control qubits, $\sigma$, in the Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubit, varies with $n$ as follows.
@@
From above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ ancilla qubits, as presented above.
@@
It is now understood that the depth of $n$-MCT can be achieved in $\mathcal{O}(\log_{2}n)$, which can be explicitly written $c\log_{2}n$, where $c$ needs to be properly estimated, for exact trade-offs which is the main motivation in this paper. In Proposition 3, we have shown that in the MCT decomposition using two clean ancilla with the technique from khattar24 , $c$ is strictly greater than 3, in fact, it is around 4. However, using the ancilla - Toffoli depth trade-off, $c$ can be further reduced to a certain extent. In the following subsection, we demonstrate that with the conditionally clean ancilla technique, $c$ always remains strictly greater than 1, regardless of the number of ancilla used.

### v2

@@ I Introduction
Quantum gates are the fundamental building blocks of quantum circuits. Unlike classical gates, quantum gates are inherently reversible and are mathematically represented by unitary matrices. The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations, reversible computing, and oracle constructions wang2024boosting ; wang2024optimal ; wang2025comprehensive ; wang2024Minimum ; wang2024Reducing ; wang2023higher . However, the Toffoli gate’s high resource demands, particularly in terms of T-Count, T-Depth, and ancilla qubits, can greatly influence the efficiency of fault-tolerant quantum circuits. Consequently, optimizing its implementation is crucial for reducing computational overhead, minimizing error rates, and enhancing scalability, thereby making it indispensable for practical large-scale quantum computations.
An $n$-controlled Toffoli is an $(n+1)$-qubit gate, where $n$ many qubits control the outcome of a single target qubit. Like ($2$-controlled) Toffoli, Multi Controlled Toffoli (MCT) also has a huge application in quantum arithmetic, error correction, and implementation of quantum algorithms. For example, in Grover’s Search, implementing the $n$-bit AND function requires an $n$-controlled Toffoli ($n$-MCT) gate. As MCT cannot be implemented in its original form (so far), it needs to be decomposed into Clifford + Toffoli, and eventually to the Clifford + T gate set. Throughout the paper, By Toffoli, we refer to $n$-controlled Toffoli with $n=2$.
It is well known that decomposing/ designing larger circuits with smaller components requires additional qubits, known as ancilla qubits, to reduce depth. The important task of optimizing the number of gates, the depth, and the number of additional qubits has received serious attention for a long time, and one may refer to the seminal work of Moore and Nilsson moore01 two decades back. A more recent and comprehensive discussion is available in jiang20 , and we like to quote from that work, “Can we characterize the relationship between the number of ancilla and the possible optimal depth?” The paper jiang20 is concerned with CNOT, whereas in this paper we concentrate on Toffoli. However, the fundamental motivation of quantum circuit synthesis revolves around this very question.
@@
Before proceeding further, let us explain the concept of conditionally clean ancilla that will be repeatedly referred to in this work. Generally, ancilla qubits are of two types, namely the clean ancilla, which is initialized to |$\ket{0}$ at the beginning of the circuit, and the dirty ancilla, which is initialized to some unknown quantum state at the beginning. The standard practice is to re-initialize the ancilla qubit to its initial state at the end of the computation. Conditionally clean ancilla nie24 ; khattar24 is the set of working (control) qubits, derived from the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation. Although earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancilla derived from the clean ones only. Very recently, the conditionally clean ancilla technique has gained significant attention, including its applications in quantum adders remaud25 .
The resource optimization metrics under consideration are the Toffoli count, Toffoli depth, and Ancilla count. The Toffoli count and Toffoli depth are further refined with the T-Count and T-Depth as one may refer to Table 1.
@@ I.1 Organization and Contributions
In Section II, we proceed with the preliminaries. Section III is a warm-up section where we first present the existing results on the Clifford $+$ T decomposition of ($2$-controlled) Toffoli and summarize the results in terms of T-Count, T-Depth, and ancilla requirements in Table 1. Additionally, we also explain the recent developments of MCT decompositions describing the existing best results in terms of Toffoli count, Toffoli depth, and ancilla, which can be subsequently decomposed into T-Count and T-Depth, as summarized in Table 2. Given that there are several developments in very recent times, this section provides a holistic view of the existing results. Based on this, we present our contributions.
In section IV, we explore the trade-off between the (clean) ancilla and the Toffoli depth using the existing techniques related to conditionally clean ones. In Section IV.1, we take a different look at viewing the MCT circuit decomposition using the conditionally clean ancilla technique of khattar24 and enumerate their exact Toffoli count. Following the trade-off, we show the reduction in Toffoli depth, and therefore in T-Depth (Construction 1, Example 1) compared to the recent work of Khattar and Gidney khattar24 , by introducing additional clean ancilla into the circuit, while keeping the Toffoli count constant. These results are shown in Section IV.2. Then, in Section IV.3, we identify the limitation of this technique in Theorem 3, showing that this direction cannot reduce the Toffoli depth to $\lceil\log_{2}n\rceil$, though it is of order $\mathcal{O}(\log_{2}n)$.
In Section V, we prove within a general framework that the exact Toffoli depth in the ($2$-controlled) Toffoli decomposition of an $n$-MCT cannot be less than $\lceil\log_{2}n\rceil$, regardless of the number of ancilla qubits used. In fact, this lower bound can be achieved for T-Depth as well, by the construction of jaques19 . That is, using the technique of jaques19 , we can obtain an $n$-MCT by further decomposing into the Clifford + T gate set with an exact T-Depth of $\lceil\log_{2}n\rceil$ too, using $22n-$ 2 ancilla, and a T-Count of $4(n-1)$. Moreover, using the logical-AND circuit by Gidney gidney18 , the ancilla count can be reduced to $n-2$ keeping the T-Count constant. However, the exact T-Depth in the case of gidney18 becomes one more, i.e., $\lceil\log_{2}n\rceil+1$. To highlight our contribution, we are looking at the exact counts and depth instead of their complexity order.
@@ II Preliminaries
However, the Clifford gates alone do not constitute a universal set of quantum gates because certain gates, for example, the T $=\sqrt{\text{S}}$ gate, cannot be arbitrarily approximated using only Clifford operations. Therefore, the Clifford group, when augmented with the T gate, forms a universal quantum gate set for quantum computation.
@@
In quantum computing, the Multi-Controlled Toffoli (MCT) gates play a crucial role in the design of complex quantum algorithms, including error correction codes and arithmetic operations. Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancillary qubits. These optimizations are essential for practical quantum computation, where resource efficiency is a critical consideration. In this regard, we present a comprehensive overview of the existing benchmarks for Clifford + T decompositions of single and multi-controlled Toffoli gates in the following section.
Let us now describe the idea of conditionally clean ancilla from khattar24 , as described in Figure 2. Given a clean ancilla, a Toffoli gate is implemented targeting the clean ancilla. If the clean ancilla is reversed, it means both the control qubits were |$\ket{1}$; thus, applying X gates on the control qubits will change them to |$\ket{0}$ and consequently can be used as conditionally clean ancilla in the next rounds. Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancilla, and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits. Once the information from all the control qubits is accumulated in a few qubits, another (smaller) MCT gate is implemented with these qubits along with the ancilla as the control qubits and the original target as its target, so that if the ancilla was not modified in the first step, the target will also not be modified.
@@
Case II: At least one of $C_{i}=\ket{0}$, for some $i\in[0,4]$. If one of $C_{0}$ or $C_{1}$ is |$\ket{0}$, then ancilla =|$\ket{0}$, and the target will not be flipped in Step III. If $C_{0}=C_{1}=\ket{1}$, but one of $C_{2},C_{3},C_{4}$ is |$\ket{0}$, then ancilla becomes |$\ket{1}$, and after applying X-gate on $C_{0},C_{1}$ will make, $C_{0}=C_{1}=\ket{0}$. If one of $C_{2},C_{3}$ is |$\ket{0}$, then $C_{0}$ and consequently, $C_{1}$ also remains at |$\ket{0}$ state. Similarly, if $C_{4}=\ket{0}$, then also $C_{1}$ remains |$\ket{0}$, and in both scenario, the target will not be flipped in Step III.
@@ III.2 Multi-Controlled Toffoli Decomposition
In Feb 2024, Nie et al. nie24 first used the notion of conditionally clean ancilla qubits derived from an existing (clean or dirty) ancilla and proposed a novel circuit decomposition (Figure 10) for the $n$-controlled Pauli gates using $\mathcal{O}(n)$ Toffoli gates. The outer layer of their MCT decomposition follows from gidney15 , and the inner layer has been parallelized to obtain an overall Toffoli depth of $\mathcal{O}(\log_{2}n)$, compared to the $\mathcal{O}(\log_{2}n)$ Toffoli depth in gidney15 . Additionally, by improving the design of a quantum incrementer, they developed an MCT circuit with a Toffoli count of $\mathcal{O}(n)$, and a Toffoli depth of $\mathcal{O}(\log^{2}n)$ without requiring any additional ancilla qubits. In this design, although the MCT decomposition does not use any additional ancilla, implementing the quantum incrementer requires one ancilla. Since our primary focus here is to reduce Toffoli depth while increasing the clean ancilla count, we ignore the zero-ancilla implementation here.
@@
The MCT circuit decomposition proposed by nie24 using a single clean ancilla has a minimum Toffoli count of $44n+$ 4 and an exact Toffoli depth of $20\log_{2}n$.
@@ IV.1 Exact Enumeration of Toffoli Depth proposed by Khattar and Gidney (2024) in a Different Lens
Using this new approach, the $5$-MCT circuit using a conditionally clean ancilla technique, presented in Figure 2, can be seen as follows.
@@
where the Toffoli depth for Step I and Step II combined is 3, which can also be verified from the actual circuit from Figure 2. Moreover, a $2$-controlled Toffoli needs to be implemented in Step III to modify the target.
@@
In the direction of enumerating the exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancilla, we have the following lemma.
@@
From the above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using 2 ancilla qubits, as proposed in khattar24 .
@@
where $k\in\mathbb{N}$, with $k\geq 2$. Since here, we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range and still show that $3\log_{2}n$ is strictly less than the corresponding depths, as follows.
@@
Although the above proposition establishes that the exact Toffoli depth using two clean ancillae, as per khattar24 , is lower bounded by $3\log_{2}n$, this bound is not tight and is closer to $4\log_{2}n$. In the following section, we analyze the trade-off between Toffoli depth and clean ancilla, demonstrating that with additional ancilla, the exact Toffoli depth can be reduced to $2\log_{2}n$.
@@ IV.2 Exact Trade-Off between Toffoli Depth and Clean Ancillae
The schematic diagram of the MCT decomposition using $m=m_{1}+m_{2}$ many ancillae is similar to the diagram we used for the exact Toffoli depth enumeration in Section IV.1.
@@
The Toffoli depth of the MCT decomposition using $m=m_{1}+m_{2}$ many ancillae can be estimated from the following lemma.
@@
The exact Toffoli depth of Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancilla following the conditionally clean ancilla technique, varies with $n$ as follows.
@@
The number of control qubits, $\sigma$, in Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubit, varies with $n$ as follows.
@@
From the above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ ancilla qubits, as presented above.
@@
It is now understood that the depth of $n$-MCT can be achieved in $\mathcal{O}(\log_{2}n)$, which can be explicitly written $c\log_{2}n$, where $c$ needs to be properly estimated, for exact trade-offs which is the main motivation in this paper. In Proposition 3, we have shown that in the MCT decomposition using two clean ancillae with the technique from khattar24 , $c$ is strictly greater than 3; in fact, it is around 4. However, using the ancilla - Toffoli depth trade-off, $c$ can be further reduced to a certain extent. In the following subsection, we demonstrate that with the conditionally clean ancilla technique, $c$ always remains strictly greater than 1, regardless of the number of ancilla used.